# PersonalVideo: High ID-Fidelity Video Customization without Dynamic and Semantic Degradation

### v1

@@ 5.3 Quantitative Results
We present the quantitative results in Tab. 1 and evaluate 1000 generated videos for 20 identities with 50 prompts from these perspectives: (1) Face Similarity: the ID cosine similarity to evaluate ID fidelity, with ID embeddings extracted using Antelopev2 [11], which is different from the face recognition models in our framework. (2) Dynamic Degree: we use VBench [23], an effective benchmark to compute video dynamics. Besides, we use well-known metrics for video evaluation. As observed, LoRA suffers from inferior face similarity due to the tuning-inference gap. MagicMe gets better face similarity yet degraded dynamics and text alignment. In contrast, our PersonalVideo significantly surpasses previous methods, especially for face similarity and dynamic degree. It achieves high ID fidelity and effectively preserves the ability of the original T2V model, which is consistent with the qualitative results. Additionally, we present the comprehensive user study in Appendix.
@@
(a) The non-reconstructive training.
@@
(b) Different steps to inject the identity.
@@
(c) Different layers to inject the identity.
Table 2: Quantitative ablation studies of our method.
5.4 Compatibility with LoRAs and ControlNet
We showcase that our framework enables controllable generation and demonstrates excellent compatibility with existing fine-grained condition modules, such as style LoRAs and ControlNet. As shown in Fig. 8, We use the Civitai community models to show that it operates effectively with these weights, even though it was not specifically trained on them. As shown in Fig. 8, the first row displays the results from the RCNZ Cartoon 3D model, while the second row and the third row highlight the outcomes from the GuoFeng RealMix and GuoFeng model. Our approach consistently delivers reliable facial preservation and effective motion generation, which has great compatibility with these customized style LoRAs. Besides, as shown in Fig. 9, we effectively utilize ControlNet with reference motion to render identities in the desired poses accurately. It highlights the robust generalization capabilities of our method, which can be seamlessly integrated with existing models.
@@ 5.4 Compatibility with LoRAs and ControlNet
5.5 Ablation Study
Non-Reconstructive Training. To validate the impact of proposed non-reconstructive training, we conduct a detailed ablation study in Fig. 10 and Tab. 2(a). They show the comparison between our PersonalVideo trained on the T2I model or T2V model, including with and without Prompt Augmentation. As observed, tuning on the T2I Model gets inferior ID fidelity due to the tuning-inference gap, which also exacerbates the distribution shift which leads to the blurred background. In contrast, tuning on the T2V model bridges the gap to achieve better ID fidelity with the preserved ability of semantic following. On the other hand, tuning without simulated prompt augmentation overfits the reference images and disrupts the original capability of the semantic following, which manifests as an inability to precisely modify the background with reduced CLIP score. With the introduction of simulated prompt augmentation, this overfitting can be significantly reduced.
Different Steps to Inject the Identity. We also conduct the ablation studies in Fig. 11 and Tab. 2(b), which illustrate the improvement in motion dynamics of our Isolated Identity Adapter to inject the identity only in the last quarter of denoising steps. As the denoising steps for injecting identity become more concentrated in the later stages, the motion dynamics of the generated videos improve accordingly. This aligns with our observations and validates the effectiveness of our design.
Different Layers to Inject the Identity. Fig. 12 and Tab. 2(c) verify the improvement in semantic following of our SIA to inject the identity only on the spatial self-attention layer. As observed, injecting only on the cross-attention layer gets inferior ID fidelity with the reference images and disrupts the original capability of semantic following, such as the losing of exquisite armor. Although injecting on both self-attention and cross-attention slightly achieves better ID fidelity, it still damages to the semantic following.
6 Conclusion & Limitation
In conclusion, we present PersonalVideo, a novel framework designed for identity-specific video generation, achieving high identity fidelity while preserving the motion dynamics and semantic following of the original T2V model. By applying direct supervision on generated videos with a learnable module, we successfully bridge the tuning-inference gap, mitigating the degradation. Furthermore, the simulated prompts augmentation enhances robustness, allowing for high-quality results even with minimal reference. Our method demonstrates superior performance over prior approaches, offering a flexible, efficient, and scalable solution for video personalization within the AIGC community.
Supplementary Material
Figure 13: User Study. Our PersonalVideo achieves the best human preference compared with other baseline methods.
Figure 14: Comparison with Textual Inversion (TI). As observed, the results of TI present low ID fidelity for 400 steps. While training for an extra 400 steps, it still suffers from the inferior ID fidelity. In comparison, our method significantly improves ID fidelity while preserving the original dynamics and semantics, which trains with non-reconstructive framework for 400 steps after initializing with a 400-step-trained TI.
Figure 15: Comparison with different inference models. For a fair comparison, we adopt the original AnimateDiff during inference. As observed, it exhibits high ID fidelity without dynamic and semantic degradation similar to AnimateLCM for the same steps, demonstrating that it does not introduce a significant gap.
@@ 6 Conclusion & Limitation
6.1 Implementation Details
Training Details. During training, we optimize the textual token with the learning rate for 1e-3 and the batch size fixed at 4. Each identity training consists of approximately 400 optimization steps. Then we learn the Isolated Identity Adapter for 400 iterations with a learning rate of 1e-4 with the batch size 1. We default to using AdamW optimizer with the default betas set to 0.9 and 0.999. The epsilon is set to the default 1e-8 and the weight decay is set to 1e-2. During inference, we use 25 steps of DDIM sampler and classifier-free guidance with a scale of 7.5 for all baselines. We generate 16-frame videos with 512 × 512 spatial resolution and 8 fps. All experiments are conducted on a single NVIDIA A800 GPU.
Baseline Details. We compare our method with both optimization methods, such as Magic-Me and LoRA, and encoder-based methods such as PuLID and ID-Animator. Specifically, Magic-Me is a recent T2V customization method that trains extended keywords on the Stable Diffusion and injects it into AnimateDiff. Besides, we compare with LoRA which also has a two-stage training. Similarly, it optimizes a keyword with Textual Inversion in the first stage and trains the LoRA layers in the second stage, which uses traditional reconstructive loss during training. For the encoder-based methods, we integrate the pre-trained encoder of PuLID, a state-of-the-art T2I customization method, with Animatediff to generate identity-specific videos. And we also compare with ID-Animator, a recent encoder-based method, which is trained on a large video dataset.
6.2 User Study
To further assess the effectiveness of our approach, we perform a human evaluation comparing our method with existing T2V identity customization techniques. We invite 12 people to review 50 sets of generated video results. For each set, we provide reference images alongside videos created using the same seed and text prompt across different methods. We evaluate the quality of the generated videos on four criteria: Identity Fidelity (the resemblance of the generated object to the reference image), Text Alignment (how well the video corresponds to the text prompt), Dynamic Degree (the dynamic degree of motion in the video), and Overall Quality (the overall satisfaction of users with the video quality). As illustrated in Fig. 13, our PersonalVideo receives significantly higher user preference across all evaluation metrics, demonstrating its effectiveness.
6.3 Discussion about the First Stage
Following the previous T2V subject customization method [45], we adopt a two-stage training strategy and use Textual Inversion to optimize the keyword, which aims to achieve a coarse identity with original motion dynamic and semantic following, as shown in Fig. 14. Then we use our non-reconstructive framework to train the model with typically 400 steps to refine the identity. Compared with Textual Inversion with 800 training steps, our method significantly achieves superior ID fidelity without introducing additional training steps.
6.4 Discussion about the Inference Model
To ensure a fair comparison with baseline methods, we use the original AnimateDiff for inference in our experiments. Although this differs from the fast sampling model AnimateLCM used during training, it does not introduce a significant gap. As shown in Fig. 15, the results of both models exhibit similar ID fidelity, with neither showing degradation in dynamics or semantics. Intuitively, this might be because AnimateLCM is a distilled version of AnimateDiff, leading to comparable performance.
6.5 More Results
As shown Fig. 16 and Fig. 17, we present more identity customization results of PersonalVideo, including few or just one reference image. They showcase it achieves high ID fidelity and preserves original motion dynamics and semantic following, which provides further evidence of its promising performance and robustness.
7 Ethics Statement
Our main objective in this work is to empower novice users to generate visual content creatively and flexibly. However, we acknowledge the potential for misuse in creating fake or harmful content with our technology. Therefore, we believe it’s essential to develop and implement tools to detect biases and malicious use cases to promote safe and equitable usage.
8 Reproducibility Statement
We make the following efforts to ensure the reproducibility of PersonalVideo: (1) Our training and inference codes together with the trained model weights will be publicly available. (2) We provide training details in the appendix (Sec. 6.1), which is easy to follow. (3) We provide the details of the human evaluation setups in the appendix (Sec. 6.2).

### v2

@@ 5.3 Quantitative Results
We present the quantitative results in Tab. 1 and evaluate 1000 generated videos for 20 identities with 50 prompts from these perspectives: (1) Face Similarity: the ID cosine similarity to evaluate ID fidelity, with ID embeddings extracted using Antelopev2 [11], which is different from the face recognition models in our framework. (2) Dynamic Degree: we use VBench [23], an effective benchmark to compute video dynamics. Besides, we use well-known metrics for video evaluation. As observed, DreamBooth suffers from inferior face similarity due to the tuning-inference gap. MagicMe gets better face similarity yet degraded dynamics and text alignment. In contrast, our PersonalVideo significantly surpasses previous methods, especially for face similarity and dynamic degree. It achieves high ID fidelity and effectively preserves the ability of the original T2V model, which is consistent with the qualitative results.
5.4 User Study
To further assess the effectiveness of our approach, we perform a human evaluation comparing our method with existing T2V identity customization techniques. We invite 15 people to review 50 sets of generated video results. For each set, we provide reference images alongside videos created using the same seed and text prompt across different methods. We evaluate the quality of the generated videos on four criteria: Identity Fidelity (the resemblance of the generated object to the reference image), Text Alignment (how well the video corresponds to the text prompt), Dynamic Degree (the dynamic degree of motion in the video), and Overall Quality (the overall satisfaction of users with the video quality). As illustrated in Fig. 7, our PersonalVideo receives significantly higher user preference across all evaluation metrics, demonstrating its effectiveness.
5.5 Compatibility with LoRAs
We also showcase that our framework demonstrates excellent compatibility with existing fine-grained condition modules, such as style LoRAs. As shown in Fig. 8, We use the Civitai community models to show that it operates effectively with these weights, even though it was not specifically trained on them. The first row displays the results from the RCNZ Cartoon 3D model [2], while the second row and the third row highlight the outcomes from the GuoFeng RealMix [1] and GuoFeng model [4]. Our approach consistently delivers reliable facial preservation and effective motion generation, which has great compatibility with these customized style LoRAs.
Figure 10: Ablation study for the Semantic Consistency Reward. As observed, tuning without SCR suffers from the dynamic degradation and the inferior semantic following, i.e., by the lake.
5.6 Ablation Study
Non-Reconstructive Training. To validate the impact of proposed non-reconstructive training, we conduct a detailed ablation study in Fig. 9 and Tab. 2. They show the comparison between our PersonalVideo trained on the T2I model or T2V model, including with and without Prompt Augmentation. As observed, tuning on the T2I Model gets inferior ID fidelity due to the tuning-inference gap, which also exacerbates the distribution shift which leads to the blurred background. In contrast, tuning on the T2V model bridges the gap to achieve better ID fidelity with the preserved ability of semantic following. On the other hand, tuning without simulated prompt augmentation overfits the reference images and disrupts the original capability of the semantic following, which manifests as an inability to precisely modify the background with reduced CLIP score. With the introduction of simulated prompt augmentation, this overfitting can be significantly reduced.
Semantic Consistency Reward. To demonstrate the effectiveness of our SCR loss, we conduct the experiments in Fig. 10 and Tab. 3. As observed, the results of tuning without SCR loss suffer from the dynamic degradation in the first example, which exhibits less movements. Besides, the results has inferior semantic following in the second example, i.e., by the lake. They verify that our proposed SCR effectively preserve the desired dynamics and semantics.
Different Steps to Inject the Identity. We also conduct the ablation studies in Fig. 11 and Tab. 4(a), which illustrate the improvement in motion dynamics of our Isolated Identity Adapter to inject the identity only in the last quarter of denoising steps. As the denoising steps for injecting identity become more concentrated in the later stages, the motion dynamics of the generated videos improve accordingly. This aligns with our observations and validates the effectiveness of our design.
6 Conclusion & Limitation
In conclusion, we present PersonalVideo, a novel framework designed for identity-specific video generation, achieving high identity fidelity while preserving the motion dynamics and semantic following of the original T2V model. By applying direct reward supervision on generated videos, we successfully bridge the tuning-inference gap, mitigating the degradation. Furthermore, the simulated prompts augmentation enhances robustness, allowing for high-quality results even with minimal reference. Our method demonstrates superior performance over prior approaches, offering a flexible, efficient, and scalable solution for video personalization within the AIGC community.
Supplementary Material
7 Implementation Details
Training Details. For AnimateDiff, we use Stable Diffusion 1.5 with Realistic Vision [3] during training and inference. During training, we learn the Isolated Identity Adapter for 800 iterations with a learning rate of 1e-4 with the batch size 1. We default to using AdamW optimizer with the default betas set to 0.9 and 0.999. The epsilon is set to the default 1e-8 and the weight decay is set to 1e-2. For Hunyuanvideo, we employ the AdamW optimizer configured with a learning rate of 2e-5 and a weight decay parameter of 1e-4 for 4000 training steps. During inference, we use 50 steps of DDIM sampler and classifier-free guidance with a scale of 7.5 for all baselines. We generate 16-frame videos with 512 × 512 spatial resolution for AnimateDiff and 61-frame videos with 720 × 1280 or 512 × 768 spatial resolution for HunyuanVideo. All experiments are conducted on a single NVIDIA A800 GPU.
Baseline Details. We compare our method with both optimization methods, such as MagicMe and Dreambooth-LoRA, and encoder-based methods such as IDAnimator and ConsisID. Specifically, Magic-Me is a recent T2V customization method that trains extended keywords on the Stable Diffusion and injects it into AnimateDiff. Besides, we compare with Dreambooth-LoRA, which uses traditional reconstructive loss during training. For a fair comparison, we train them for the same steps with PersonalVideo. For the encoder-based methods, we compare with ID-Animator and ConsisID, the recent encoder-based methods, which are both trained on the large video dataset.
8 More Comparison
We provide more comparison including more base models and different number of the references in Fig. 12, Fig. 13, and Fig. 14. As shown, both Dreambooth and MagicMe suffer from inferior ID fidelity. Besides, MagicMe has a severe misalignment of the prompt, e.g., tuning head. In contrast, our PersonalVideo maintains higher ID fidelity without dynamic and semantic degradation, which is consistent with results on the DiT-based model.
9 More ablation study
Fig. 15 and Tab. 4(b) verify the improvement in semantic following of our Isolated Identity Adapter to inject the identity only on the spatial self-attention layer. As observed, injecting only on the cross-attention layer gets inferior ID fidelity with the reference images and disrupts the original capability of semantic following, such as the losing of exquisite armor. Although injecting on both self-attention and cross-attention slightly achieves better ID fidelity, it still damages to the semantic following.
@@ 9 More ablation study
Table 2: Quantitative ablation of the non-reconstructive training.
Face ($\uparrow$) CLIP-T ($\uparrow$) Dynamic ($\uparrow$)
w/o SCR 61.08 26.38 13.22
w/ SCR 61.05 28.59 17.85
Table 3: Quantitative ablation of Semantic Consistency Reward.
@@
(a) Different steps to inject the identity.
@@
(b) Different layers to inject the identity in the Unet-based model.
Table 4: Quantitative ablation studies of Isolated Identity Adapter.
Figure 12: Qualitative comparison on Animatediff. As observed, both Dreambooth and MagicMe suffer from inferior ID fidelity. Besides, MagicMe has a semantic following degradation, e.g., tuning head. In contrast, our PersonalVideo maintains high ID fidelity and preserve the original motion dynamics and semantic following, significantly surpassing others.
Figure 13: More comparisons on HunyuanVideo. As observed, Dreambooth suffers from inferior ID fidelity, while our PersonalVideo maintains higher ID fidelity without dynamic and semantic degradation, which is consistent with Fig. 12.
Figure 14: More comparison for a single reference. While ConsisID and Dreambooth suffer from the inferior ID fidelity, as well as severe degradation of motion dynamics and semantic following, e.g. the stool by the lake, our PersonalVideo achieves robust customization with high ID fidelity and preserved motion dynamics and semantic following.
10 More Results
As shown in Fig. 20, Fig. 21, Fig. 22, Fig. 23, and Fig. 24, we present more customization results of PersonalVideo, including few or just one reference image. They showcase it achieves high ID fidelity and preserves original motion dynamics and semantic following, which provides further evidence of its promising performance and robustness.
11 Reproducibility Statement
We make the following efforts to ensure the reproducibility of PersonalVideo: (1) Our training and inference codes together with the trained model weights will be publicly available. (2) We provide training details in the appendix (Sec. 7), which is easy to follow. (3) We provide the details of the human evaluation setups in the appendix (Sec. 5.4).
12 Impact Statement
Our main objective in this work is to empower novice users to generate visual content creatively and flexibly. However, we acknowledge the potential for misuse in creating fake or harmful content with our method. Thus, we believe it’s essential to develop and implement tools to detect biases and malicious use cases to promote safe and equitable usage.
@@ 12 Impact Statement
Figure 16: More results of PersonalVideo.
Figure 17: More results of PersonalVideo.
Figure 18: More results of PersonalVideo.
Figure 19: More results of PersonalVideo.
Figure 20: More results of PersonalVideo.
Figure 21: More results of PersonalVideo.
Figure 22: More results of PersonalVideo.